The capital gains yield equals which one of the following?

A. Total yield
B. Required rate of return C. Market rate of return
D. Dividend yield
E. Market price growth rate
Answer #1

option 'E' is correct

Market Price growth rate

Capital Gain yield = (P1 - P0) *100 / P0

P1 = price at end of Year 1

P0 = price at beginning of year

Capital gains yield equals to the Market price growth rate.

example Price at end of the year = $105. Price at time of purchase = $100

Capital gains yield = ($105 - $100) / $100 = 5%

So, P1 can be calculated by = P0 * (1 + g) = $100 * (1.05) = $105